The tip of a blade of an electric fan is 1.5 feet from the axis of rotation . if the fan spins at a full rate of 1,160 revolutions per minute, how many miles will a point at the tip of a blade travel in one hour ? (1 mile = 5,280 feet)
a 30pi b 40 pi c 45 pi d 48 pi e 60 pi

Answer by CharlesG2(834)

The tip of a blade of an electric fan is 1.5 feet from the axis of rotation . if the fan spins at a full rate of 1,160 revolutions per minute, how many miles will a point at the tip of a blade travel in one hour ? (1 mile = 5,280 feet)
a 30pi b 40 pi c 45 pi d 48 pi e 60 pi
C circumference = 2piR
R radius = 1.5 ft
C = 2 * pi * 1.5 = 3pi ft
1160 * rev/min * 3pi * ft/rev = 1160 * 3pi = 3480pi feet/minute
3480pi * feet/minute * 1/5280 * mile/feet = (3480/5280)pi miles/minute
3480/5280 = 348/528 = 174/264 = 87/132
(87/132)pi * miles/minute * 60 * minutes/hour = (87/132) * 60 * pi miles/hr
approx 39.55 pi miles/hr
so b. 40pi